OpenStudy (anonymous):

If you stand on a ship in a calm sea, then your height x (in ft) above sea level is related to the farthest distance y (in mi) that you can see by the equation: y=sqrt(1.8x+(x/5280)^2) (b) How high up do you have to be to be able to see 5 mi? (Round your answer to the nearest whole number.)

OpenStudy (ash2326):

our height x (in ft) above sea level is related to the farthest distance y (in mi) that you can see by the equation: \[y=\sqrt{1.8x+{(\frac{x}{5280})}^2}\] here we have to find x for which we are able to see 5 miles, so put y=5 \[\large 5=\sqrt{1.8x+{(\frac{x}{5280})}^2}\] square both sides and solve the quadratic for x, can you do it @Brent0423

OpenStudy (anonymous):

ive already tried that and i get 1.8x+x^2+696960000

OpenStudy (ash2326):

use the quadratic formula to solve for x \[ax^2+bx+c=0\] \[\large x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

-1.8+-sqrt((-1.8)^2-4(1)(696960000))/2(1)

OpenStudy (anonymous):

im getting an imaginary number, this hw is due in 16 mins can u please just help me

OpenStudy (ash2326):

let me check

OpenStudy (ash2326):

\[25=1.8x+(\frac{x}{5280})^2\] \[\frac{x^2}{5280^2}+1.8x-25=0\] Use quadratic formula \[\Large x=\frac{-1.8\pm\sqrt{(1.8)^2-4\times \frac{1}{5280^2}\times (-25)}}{2\times \frac{1}{5280^2}}\] now it's just simplification \[\Large x=\frac{-1.8\pm\sqrt{(1.8)^2+4\times \frac{1}{5280^2}\times 25}}{2\times \frac{1}{5280^2}}\]' could you solve this from here?

OpenStudy (anonymous):

i dont have calculator please tell me the answer

OpenStudy (anonymous):

hurry only 3 mins remaining!

OpenStudy (anonymous):

help!!!!!!

OpenStudy (ash2326):

we'll get \[\Large x=\frac{-1.8\pm \sqrt{3.24000358701}}{7.17401286\times 10^{-8}}\] \[\Large x=\frac{-1.8\pm 1.80000099639}{7.17401286\times 10^{-8}}\] we'll get two x \[x=13.889, -5.00\times 10^{17}\] the positive x is your answer

OpenStudy (ash2326):

Did you understand @Brent0423 ?